The probability that Tope passes an exam is 1/3 and probabilities that his brother

and sister pass the same exam are 2/5 and 3/8 respectively. Find the probability that

the three of them will fail the exam.


1
Expert's answer
2022-04-16T04:13:01-0400

Let events "A, B, C -" Tope, his brother and his sister pass an exam.

"P(A) =\cfrac{1} {3} ,P(B) =\cfrac{2} {5} ,P(A) =\cfrac{3} {8}."

Events that they fail an exam are "\bar{A}, \bar{B}, \bar{C}" respectively, their probabilities

"P(\bar{A} ) =1-\cfrac{1} {3}=\cfrac{2} {3} ,\\
P(\bar{B} ) =1-\cfrac{2} {5}=\cfrac{3} {5} ,\\
P(\bar{C} ) =1-\cfrac{3} {8}=\cfrac{5} {8}."


The events "\bar{A}, \bar{B}, \bar{C}" are independent,

"P(\bar{A}\cap\bar{B}\cap\bar{C}) =P(\bar{A}) \cdot P(\bar{B}) \cdot P(\bar{C}) =\\
=\cfrac{2} {3} \cdot\cfrac{3} {5} \cdot\cfrac{5} {8} =\cfrac{1} {4} ."